# Hybrid Feedback for Three-dimensional Convex Obstacle Avoidance (Extended version)

### v3

@@ IX-C Proof of Lemma 3
Hence, according to [23, Proposition 6.10], since (36) holds for all $\xi\in\mathcal{F}\setminus\mathcal{J}$, there exists a nontrivial solution to $\mathcal{H}$ for each initial condition in $\mathcal{K}$. Finite escape time can only occur through flow. They can neither occur for $\mathbf{x}$ in the set $\mathcal{F}_{1}^{\mathcal{W}}$, as this set is bounded as per definition (23), nor for $\mathbf{x}$ in the set $\mathcal{F}_{0}^{\mathcal{W}}$ as this would make $\mathbf{x}^{\intercal}\mathbf{x}$ grow unbounded, and would contradict the fact that $\frac{d}{dt}(\mathbf{x}^{\intercal}\mathbf{x})\leq 0$ in view of the definition of $\mathbf{u}(\mathbf{x},\mathbf{h},\mathbf{a},0,s)$. Therefore, all maximal solutions do not have finite escape times. Furthermore, according to (31), $\mathbf{x}^{+}=\mathbf{x},$ and from the definition of the update law in (27) and (30), it follows immediately that $\mathbf{J}(\mathcal{J})\subset\mathcal{K}$. Hence, the solutions to the hybrid closed-loop system (31) cannot leave $\mathcal{K}$ through jump and, as per [23, Proposition 6.10], all maximal solutions are complete.
@@ IX-D Proof of Theorem 1
Forward invariance and stability: The forward invariance of the obstacle-free set $\mathcal{K}$, for the hybrid closed-loop system (31), is immediate from Lemma 3. We next prove the stability of $\mathcal{A}$ using [23, Definition 7.1].
@@
Consequently, for every $\rho>0$, one can choose $\sigma\in(0,\min\{\mu_{1},\mu_{2},\mu_{3},\rho\})$ such that for all initial conditions $\xi(0,0)$ with $d(\xi(0,0),\mathcal{A})\leq\sigma$, one has $d(\xi(t,j),\mathcal{A})\leq\rho$ for all $(t,j)\in\text{ dom }\xi,$ where $d(\xi,\mathcal{A})^{2}=\underset{(\mathbf{0},\bar{\mathbf{h}},\bar{\mathbf{a}},\bar{m},s)\in\mathcal{A}}{\inf}(\lVert\mathbf{x}\rVert^{2}+\lVert\mathbf{h}-\bar{\mathbf{h}}\rVert^{2}+\lVert\mathbf{a}-\bar{\mathbf{a}}\rVert^{2}+(m-\bar{m})^{2}+(s-\bar{s})^{2})=\lVert\mathbf{x}\rVert^{2}.$ Hence, according to [25, Definition 3.1], the target set $\mathcal{A}$ is stable for the hybrid closed-loop system (31). Next, we proceed to establish the convergence properties of the set $\mathcal{A}$.
Attractivity: We aim to show that for the proposed hybrid closed-loop system (31), the target set $\mathcal{A}$ is globally attractive in the set $\mathcal{K}$ using [25, Defintion 3.1 and Remark 3.5]. In other words, we prove that for all initial conditions $\xi(0,0)\in\overline{\mathcal{F}}\cup\mathcal{J}=\mathcal{K}$, every maximal solution $\xi$ to the hybrid closed-loop system is complete and satisfies
@@
Hence, the target set $\mathcal{A}$ is globally attractive in the set $\mathcal{K}$ for the proposed hybrid closed-loop system (31). In addition, since the set $\mathcal{A}$ is stable for the hybrid closed-loop system (31), it is globally asymptotically stable in the set $\mathcal{K}$ for the hybrid closed-loop system (31) as per [25, Remark 3.5].
@@ IX-E Proof of Lemma 4
This proof is by contradiction. First, note that for any $\mathbf{q}\in\mathbb{S}^{2}$, the matrix $\mathbf{P}(\mathbf{q})$ is positive semidefinite. Therefore, for any vector $\mathbf{s}\in\mathbb{R}^{n}$, one has $\mathbf{s}^{\top}\mathbf{P}(\mathbf{q})\mathbf{s}\geq 0$. Let us assume that there exists $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{p},\mathbf{q})$ such that $\mathbf{x}_{\pi}^{\top}\mathbf{P}(\mathbf{q})\mathbf{x}_{\pi}=0$. Since $\mathbf{x}_{\pi}\neq\mathbf{0}$, one has $\mathbf{P}(\mathbf{q})\mathbf{x}_{\pi}=\mathbf{0}$. Therefore, the vector $\mathbf{x}_{\pi}$ is normal to the hyperplane $\mathcal{P}(\mathbf{p},\mathbf{q})$. This implies that the plane $\mathcal{P}(\mathbf{p},\mathbf{q})$ is a supporting hyperplane [26, Section 2.5.2] to the convex set $\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})$ at $\mathbf{x}$, where $\beta=d(\mathbf{x},\mathcal{O}_{i})-r_{a}\in[0,\gamma]$. Therefore, the set $\mathcal{D}_{r_{a}}^{\circ}(\mathcal{O}_{i})\cap\mathcal{P}(\mathbf{p},\mathbf{q})$ is an empty set.

### v4

@@ IX-C Proof sketch for Lemma 3
According to [20, Proposition 6.10], since (35) holds for all $\xi\in\mathcal{F}\setminus\mathcal{J}$, there exists a maximal solution to (31) for each initial condition in $\mathcal{K}$. Finite escape time can only occur through flow, but it cannot occur for $\mathbf{x}$ in the set $\mathcal{F}_{1}^{\mathcal{W}}$, as this set is bounded by definition (23), nor for $\mathbf{x}$ in the set $\mathcal{F}_{0}^{\mathcal{W}}$, as this would make $\mathbf{x}^{\intercal}\mathbf{x}$ grow unbounded, contradicting the fact that $\frac{d}{dt}(\mathbf{x}^{\intercal}\mathbf{x})\leq 0$ according to the definition of $\mathbf{u}(\mathbf{x},\mathbf{h},\mathbf{a},0,s)$. Therefore, all maximal solutions do not have finite escape times, and Claim (b) in [20, Proposition 6.10] does not hold.
According to (20) and (24), by construction, we have $\mathcal{F}_{m}\cup\mathcal{J}_{m}=\mathcal{K}$ for each $m\in\mathbb{M}$. Therefore, using (27), (30) and (31), it can be verified that $\mathbf{L}(\xi)\in\mathcal{F}\cup\mathcal{J}=\mathcal{K}$ for every $\xi\in\mathcal{K}$. As a result, Claim (c) in [20, Proposition 6.10] does not apply. Therefore, Claim (a) in [20, Proposition 6.10] holds. In other words, all maximal solutions to the hybrid closed-loop system (31) are complete, implying the forward invariance of the obstacle-free set $\mathcal{K}$.
@@ IX-D Proof of Theorem 1
Forward invariance and stability: The forward invariance of the obstacle-free set $\mathcal{K}$, for the hybrid closed-loop system (31), is immediate from Lemma 3. We next prove the stability of $\mathcal{A}$ using [20, Definition 7.1].
@@
Consequently, for every $\rho>0$, one can choose $\sigma\in(0,\min\{\mu_{1},\mu_{2},\mu_{3},\rho\})$ such that for all initial conditions $\xi(0,0)$ with $d(\xi(0,0),\mathcal{A})\leq\sigma$, one has $d(\xi(t,j),\mathcal{A})\leq\rho$ for all $(t,j)\in\text{ dom }\xi,$ where $d(\xi,\mathcal{A})^{2}=\underset{(\mathbf{0},\bar{\mathbf{h}},\bar{\mathbf{a}},\bar{m},s)\in\mathcal{A}}{\inf}(\lVert\mathbf{x}\rVert^{2}+\lVert\mathbf{h}-\bar{\mathbf{h}}\rVert^{2}+\lVert\mathbf{a}-\bar{\mathbf{a}}\rVert^{2}+(m-\bar{m})^{2}+(s-\bar{s})^{2})=\lVert\mathbf{x}\rVert^{2}.$ Hence, according to [22, Definition 3.1], the target set $\mathcal{A}$ is stable for the hybrid closed-loop system (31). Next, we proceed to establish the convergence properties of the set $\mathcal{A}$.
Attractivity: We aim to show that for the proposed hybrid closed-loop system (31), the target set $\mathcal{A}$ is globally attractive in the set $\mathcal{K}$ using [22, Defintion 3.1 and Remark 3.5]. In other words, we prove that for all initial conditions $\xi(0,0)\in\overline{\mathcal{F}}\cup\mathcal{J}=\mathcal{K}$, every maximal solution $\xi$ to the hybrid closed-loop system is complete and satisfies
@@
Hence, the target set $\mathcal{A}$ is globally attractive in the set $\mathcal{K}$ for the proposed hybrid closed-loop system (31). In addition, since the set $\mathcal{A}$ is stable for the hybrid closed-loop system (31), it is globally asymptotically stable in the set $\mathcal{K}$ for the hybrid closed-loop system (31) as per [22, Remark 3.5].
@@ IX-E Proof of Lemma 4
This proof is by contradiction. First, note that for any $\mathbf{q}\in\mathbb{S}^{2}$, the matrix $\mathbf{P}(\mathbf{q})$ is positive semidefinite. Therefore, for any vector $\mathbf{s}\in\mathbb{R}^{n}$, one has $\mathbf{s}^{\top}\mathbf{P}(\mathbf{q})\mathbf{s}\geq 0$. Let us assume that there exists $\mathbf{x}\in\mathcal{N}_{\gamma}(\mathcal{D}_{r_{a}}(\mathcal{O}_{i}))\cap\mathcal{P}(\mathbf{p},\mathbf{q})$ such that $\mathbf{x}_{\pi}^{\top}\mathbf{P}(\mathbf{q})\mathbf{x}_{\pi}=0$. Since $\mathbf{x}_{\pi}\neq\mathbf{0}$, one has $\mathbf{P}(\mathbf{q})\mathbf{x}_{\pi}=\mathbf{0}$. Therefore, the vector $\mathbf{x}_{\pi}$ is normal to the hyperplane $\mathcal{P}(\mathbf{p},\mathbf{q})$. This implies that the plane $\mathcal{P}(\mathbf{p},\mathbf{q})$ is a supporting hyperplane [23, Section 2.5.2] to the convex set $\mathcal{D}_{r_{a}+\beta}(\mathcal{O}_{i})$ at $\mathbf{x}$, where $\beta=d(\mathbf{x},\mathcal{O}_{i})-r_{a}\in[0,\gamma]$. Therefore, the set $\mathcal{D}_{r_{a}}^{\circ}(\mathcal{O}_{i})\cap\mathcal{P}(\mathbf{p},\mathbf{q})$ is an empty set.